# Multiple Imputation of Hierarchical Nonlinear Time Series Data with an Application to School Enrollment Data

### v1

@@ 5.2.1 Analysis Model
We are interested in estimating the relationship between NER and the Total Fertility Rate (TFR), which is a period measure of the expected number of children a woman would bear in her lifetime if she were to experience the period-specific fertility rates at each age and if she lived through the reproductive age range of 15–49. There is a well-established negative association between education and fertility in the high-fertility setting (Hirschman, 1994). One mechanism through which education is posited to have a negative effect on fertility is through the educational enrollment of children, where increased enrollment increases the cost of raising children (Axinn and Barber, 2001). Children who are enrolled in school have reduced capacity for work and may incur increased costs for caregivers through fees related to tuition, uniforms, and textbooks (Easterlin and Crimmins, 1985; Caldwell, 1982; Caldwell et al., 1985). To estimate this relationship, we use annual estimates of TFR from the United Nations World Population Prospects 2022 (United Nations (2022)). We restrict our analyses to the high-fertility context, defined here as the years where a country has TFR $>2.5$ children per woman.
@@
For each replication of each experiment, the analysis model is estimated using only the country-years that were simulated as missing and where TFR $>2.5$. We also conducted analysis model validation using a random intercept model for TFR on NER, results of which are available in the Appendix.
@@ 5.2.2 Model Implementation
GER is the auxiliary variable $\mathbf{X}$ and NER is the variable of interest $\mathbf{Y}$. The bounds of the model for $\mathbf{X}$ were set as $X_{low}=0$ and $X_{up}=\infty$, while the bounds of the model for $\mathbf{Y}|\mathbf{X}$ were set as $Y_{low}=0$ and $Y_{up}=\text{min}(X_{c,t},100)$. Imputations were created using MINTS for each replication of each experiment by running 10 chains of the MCMC algorithm until convergence was reached in the estimation phase. The total number of iterations differed across experiments, with the experiments with larger rates of simulated missingness requiring a larger number of iterations to achieve convergence. After burn-in, the number of iterations per chain ranged from 5000 to 35000. In the imputation phase, an additional 4000 iterations was run for each chain and four iterations were selected from each chain to produce a total of $M=40$ completed data sets following the procedure described in Section 3.3.2.
Imputations from the MICE and Amelia methods were created using the mice and Amelia packages in R. We set the number of imputations to $M=40$, specified the form of the imputation models as described in Section 4.4, and specified bounds of $\mathbf{X}\in[0,\infty)$ and $\mathbf{Y}\in[0,100]$, but otherwise used the default software settings.
@@ 5.2.3 Results
Table 2 summarizes the results of the analysis model validation for estimation of $Q=\beta_{1}$, the regression coefficient on NER in the linear regression of TFR on NER. MINTS results in the best overall performance, with the smallest MAE in all experiments except MAR 80%, where MINTS has the second smallest MAE after pan Fixed Effects. MINTS also has the smallest FMI in all experiments and close to nominal coverage in all but the MNAR 40% experiment. Out of the previously existing methods, pan Fixed Effects performs the best overall, with the smallest MAE in the MAR 80% experiment and the second smallest MAE in all other experiments. Pan Fixed Effects also results in good coverage for the MCAR and MAR experiments and has the closest to nominal coverage for the MNAR 40% experiment.
@@
Simulated Method MCAR MAR MNAR Missingness Rate MAE Cvg FMI MAE Cvg FMI MAE Cvg FMI 10% MICE PMM 0.61 90 20.9 0.86 71 26.3 7.29 0 28.2 pan Fixed 0.08 100 7.2 0.11 100 6.8 2.69 0 29.9 pan Random 0.07 100 5.3 0.12 100 7.8 4.08 2 88.4 Amelia TS 0.29 100 24.2 0.51 99 26.7 7.16 0 30.6 Amelia CS 0.21 100 11.4 0.40 99 11.9 6.03 0 58.9 Amelia TSCS 0.09 100 8.3 0.16 100 12.9 6.42 0 69.2 MINTS 0.04 100 2.2 0.08 100 2.4 0.42 100 10.8 40% MICE PMM 1.82 0 35.0 1.94 0 38.1 5.22 0 43.0 pan Fixed 0.05 100 13.4 0.08 100 12.8 1.02 76 74.9 pan Random 0.13 99 21.0 0.20 100 31.4 4.73 0 93.7 Amelia TS 1.34 0 32.4 1.48 0 35.5 4.77 0 40.5 Amelia CS 0.69 3 24.6 0.86 0 24.6 6.04 0 69.1 Amelia TSCS 0.22 99 22.4 0.43 74 32.4 6.42 0 71.1 MINTS 0.04 100 4.6 0.07 100 5.4 0.63 13 31.9 80% MICE PMM 2.96 0 52.1 3.08 0 54.2 pan Fixed 0.07 100 41.4 0.09 100 45.9 pan Random 1.20 0 69.3 1.29 0 76.6 Amelia TS 2.55 0 44.6 2.64 0 47.0 Amelia CS 1.45 0 57.8 1.52 0 57.3 Amelia TSCS 1.01 2 62.4 1.57 0 64.3 MINTS 0.06 100 21.8 0.11 100 25.5
We additionally considered an analysis model validation exercise using a random intercept analysis model for the regression of TFR on NER. For the 10% rate of simulated missingness, MINTS has the best performance for estimation of the fixed effect coefficient on NER in the random intercept analysis model. The pan Random Effects method tends to perform the best for estimation of the fixed effect coefficient at higher rates of missingness, while MINTS performs the best for estimation of the variance of the random intercepts. Full results for the random intercept analysis model validation exercise can be found in the Appendix.
@@ 5.3 Out-of-Sample Validation
Next, we evaluated the predictive performance of each multiple imputation method for imputing the individual values of NER that were simulated as missing. In each experiment, the performance of point estimates was evaluated using the mean absolute error, where the mean was taken over all country-years in the testing set. The performance of the predictive intervals was evaluated by checking the average interval widths and coverage of the 95% intervals with respect to the observations in the testing set, where coverage is calculated as the proportion of the intervals that contained the true left-out values of NER. We also evaluated how well each imputation method balances the trade-off between interval width and coverage for the 95% intervals using the average negatively-oriented interval score from Gneiting and Raftery (2007). For a 95% prediction interval, the interval score is defined as
@@
The out-of-sample validation evaluates the posterior predictive distribution of the missing values given the observed values. Samples from the posterior predictive distribution of the missing values from the MINTS method were obtained using a slight modification of the imputation procedure, details of which can be found in the Appendix. Details of how samples from the posterior predictive distributions were obtained for the MICE and Amelia methods can also be found in the Appendix. Medians from the estimated posterior predictive distributions were used as point estimates for the imputed values and the 0.025 and 0.975 quantiles of the estimated posterior predictive distributions were used as 95% interval estimates.
@@ 5.3.1 Results
MICE PMM performs the worst overall, with comparatively large MAE and undercoverage in all experiments. This is perhaps unsurprising, as MICE PMM is the only method that does not account for the hierarchical structure of the data. Amelia TS also suffers from similarly large MAE as MICE PMM, but manages to retain close to nominal coverage in experiments under MCAR and MAR thanks to its wider intervals. The previously existing methods that have the best performance are pan Random Effects and Amelia TSCS. These two methods perform similarly in terms of MAE within experiments, with Amelia TSCS having larger MAE overall. Amelia TSCS has narrower intervals than pan Random Effects in most experiments, where Amelia TSCS tends towards undercoverage while pan Random Effects tends towards overcoverage. Despite performing well overall, both these methods lead to larger MAE and interval scores than MINTS.
@@
We illustrate the out-of-sample validation results for the MINTS method for selected experiments in Figures 3 and 4. Observations in the training set are shown as solid circles in black for GER and red for NER. Posterior medians for the imputed values of NER are shown as red open circles, while the red shaded regions represent the 95% posterior predictive intervals. Imputed values are plotted for both the country-years that started as missing in the enrollment data set and the country-years that were simulated as missing for the testing set. The true values of NER for the country-years in the testing set are shown as solid blue diamonds. Figure 3 shows the out-of-sample validation results for Afghanistan, Belgium, Spain, and Nigeria for the MCAR 40% experiment. Overall, the posterior predictive distributions for imputed values of NER result in plausible time series trends for each country, with the majority of the observations of NER that were simulated as missing captured within the 95% intervals. Results for the same example countries are shown in Figure 4 for the MAR 80% experiment. The predictive intervals are much wider overall for the MAR 80% experiment compared to the MCAR 40% experiment. This is especially apparent for Spain, which has many observations of NER and GER in early years and thus had a high proportion of observed values that were simulated as missing in the MAR 80% experiment.
@@ 5.4 Full Enrollment Data Results
We used MINTS to create 40 multiple imputations for the missing country-years in the original enrollment data set without simulating any additional country-years as missing. Figure 5 shows the multiple imputations for Afghanistan, Belgium, Spain, and Nigeria for NER as translucent red circles along with the observed values of NER and GER from the original data set in solid red and black circles, respectively. A CSV of 40 completed data sets for NER and GER created using the MINTS method and a CSV of the median imputed values for NER and GER are provided as ancillary files.
@@
Figure 5: Results of $M=40$ imputations for NER for selected countries. Solid black and red circles indicate observed values of GER and NER, respectively, from the enrollment data set. Translucent open red circles indicate imputed values of NER, where a total of 40 imputations were created for each missing value.

### v2

@@ 5.2.1 Analysis Model
We are interested in estimating the relationship between NER and the Total Fertility Rate (TFR), which is a period measure of the expected number of children a woman would bear in her lifetime if she were to experience the period-specific fertility rates at each age and if she lived through the reproductive age range of 15–49. There is a well-established negative association between education and fertility in the high-fertility setting (Hirschman (1994)). One mechanism through which education is posited to have a negative effect on fertility is through the educational enrollment of children, where increased enrollment increases the cost of raising children (Axinn and Barber (2001)). Children who are enrolled in school have reduced capacity for work and may incur increased costs for caregivers through fees related to tuition, uniforms, and textbooks (Easterlin and Crimmins (1985); Caldwell (1982); Caldwell, Reddy and Caldwell (1985)). To estimate this relationship, we use annual estimates of TFR from the 2022 revision of the United Nations World Population Prospects (United Nations (2022)). We restrict our analyses to the high-fertility context, defined here as the years where a country has TFR $>2.5$ children per woman.
@@
For each replication of each experiment, the analysis model is estimated using only the country-years that were simulated as missing and where TFR $>2.5$. We also conducted analysis model validation using a random intercept model for TFR on NER, results of which are available in Section 4 of the Supplementary Material (Liu and Raftery (2025)).
In this validation exercise, the analysis model is uncongenial to all imputation models considered due to the omission of the analysis model outcome variable $\mathbf{Z}=\text{TFR}$ from the imputation models. This reflect our motivating setting where analysis of the multiply imputed values occurs independently from multiple imputation, and information about the outcome variable is not known during the imputation task. For comparison purposes, we also consider a data augmentation algorithm based on MINTS to represent the setting where imputation of missing values and estimation of the substantive analysis model can be conducted simultaneously.
@@ 5.2.2 Model Implementation
[…]
Multiple imputations were created using the models based on the MICE and Amelia methods following the same implementation as described in Section 4.4 with two exceptions. First, the school enrollment data was not transformed prior to using the MICE and Amelia methods, as we were unable to find a transformation that substantially improved linearity for the relationship between $\mathbf{X}$ and $\mathbf{Y}$. Second, the bounds of each model were specified as $\mathbf{X}\in[0,\infty)$ and $\mathbf{Y}\in[0,100]$. The data augmentation method based on MINTS was also implemented as described in Section 4.4.
@@ 5.2.3 Results
Table 2 summarizes the results of the analysis model validation for estimation of $Q=\beta_{1}$, the regression coefficient on NER in the linear regression of TFR on NER. MINTS has the smallest MAE out of the multiple imputation methods considered in all but the MAR 80% experiment. MINTS also has the smallest FMI in all experiments and close to nominal coverage in all but the MNAR 40% experiment. Out of the previously existing methods, pan Fixed Effects performs the best overall with the smallest MAE in the MAR 80% experiment and the closest to nominal coverage in the MNAR 40% experiment. The data augmentation approach based on MINTS has the smallest MAE and closest to nominal coverage in the MNAR experiments, but falls behind several multiple imputation methods in the experiments where the MAR assumption is satisfied.
@@
Simulated Method MCAR MAR MNAR Missingness Rate MAE Cvg FMI MAE Cvg FMI MAE Cvg FMI 10% MICE PMM 0.61 90 20.9 0.86 71 26.3 7.29 0 28.2 pan Fixed 0.08 100 7.2 0.11 100 6.8 2.69 0 29.9 pan Random 0.07 100 5.3 0.12 100 7.8 4.08 2 88.4 Amelia TS 0.29 100 24.2 0.51 99 26.7 7.16 0 30.6 Amelia CS 0.21 100 11.4 0.40 99 11.9 6.03 0 58.9 Amelia TSCS 0.09 100 8.3 0.16 100 12.9 6.42 0 69.2 MINTS 0.04 100 2.2 0.08 100 2.4 0.42 100 10.8 MINTS DA 0.28 100 0.23 100 0.34 100 40% MICE PMM 1.82 0 35.0 1.94 0 38.1 5.22 0 43.0 pan Fixed 0.05 100 13.4 0.08 100 12.8 1.02 76 74.9 pan Random 0.13 99 21.0 0.20 100 31.4 4.73 0 93.7 Amelia TS 1.34 0 32.4 1.48 0 35.5 4.77 0 40.5 Amelia CS 0.69 3 24.6 0.86 0 24.6 6.04 0 69.1 Amelia TSCS 0.22 99 22.4 0.43 74 32.4 6.42 0 71.1 MINTS 0.04 100 4.6 0.07 100 5.4 0.63 13 31.9 MINTS DA 0.27 100 0.28 100 0.53 100 80% MICE PMM 2.96 0 52.1 3.08 0 54.2 pan Fixed 0.07 100 41.4 0.09 100 45.9 pan Random 1.20 0 69.3 1.29 0 76.6 Amelia TS 2.55 0 44.6 2.64 0 47.0 Amelia CS 1.45 0 57.8 1.52 0 57.3 Amelia TSCS 1.01 2 62.4 1.57 0 64.3 MINTS 0.06 100 21.8 0.11 100 25.5 MINTS DA 0.53 100 0.54 100
We also considered a validation exercise using a random intercept analysis model for the regression of TFR on NER. For the 10% rate of simulated missingness, MINTS has the best performance for estimation of the fixed effect coefficient on NER in the random intercept analysis model. The pan Random Effects model tends to perform the best for estimation of the fixed effect coefficient at higher rates of missingness, while MINTS performs the best for estimation of the variance of the random intercepts. Full details of this analysis model validation exercise can be found in Section 4 of the Supplementary Material (Liu and Raftery (2025)).
@@ 5.3 Out-of-Sample Validation
Next, we evaluated the predictive performance of each method for imputing the individual values of NER that were simulated as missing. In each experiment, the performance of point estimates was evaluated using the mean absolute error, where the mean was taken over all country-years in the testing set. The performance of the predictive intervals was evaluated by checking the average interval widths and coverage of the 95% intervals with respect to the observations in the testing set, where coverage is calculated as the proportion of the intervals that contained the true left-out values of NER. We also evaluated how well each imputation method balances the trade-off between interval width and coverage for the 95% intervals using the average negatively-oriented interval score from Gneiting and Raftery (2007). For a 95% prediction interval, the interval score is defined as
@@
The out-of-sample validation exercise evaluates the posterior predictive distribution of the missing values given the observed values. Samples from the posterior predictive distribution of the missing values from the MINTS method were obtained using a slight modification of the imputation procedure, details of which can be found in Section 7 of the Supplementary Material (Liu and Raftery (2025)). Details of how samples from the posterior predictive distributions were obtained for the MICE and Amelia methods can also be found in Section 7 of the Supplementary Material (Liu and Raftery (2025)). Medians from the estimated posterior predictive distributions were used as point estimates for the imputed values and the 0.025 and 0.975 quantiles of the estimated posterior predictive distributions were used as 95% interval estimates.
@@ 5.3.1 Results
MICE PMM performs the worst overall, with comparatively large MAE and undercoverage in all experiments. This is perhaps unsurprising, as MICE PMM is the only method that does not account for the hierarchical structure of the data. Amelia TS suffers from similarly large MAE as MICE PMM, but retains close to nominal coverage in experiments under MCAR and MAR thanks to its wider intervals. Out of the previously existing methods, pan Random Effects and Amelia TSCS perform the best overall. Amelia TSCS tends towards undercoverage, having larger MAE and narrower intervals than pan Random Effects in most experiments, while pan Random Effects tends towards overcoverage.
@@
We illustrate the out-of-sample validation results for the MINTS method for selected experiments in Figures 3 and 4. Observations in the training set are shown as solid circles in black for GER and red for NER. Posterior medians for the imputed values of NER are shown as red open circles, while the red shaded regions represent the 95% posterior predictive intervals. Imputed values are plotted for both the country-years that started as missing in the school enrollment data set and the country-years that were simulated as missing for the testing set. The true values of NER for the country-years in the testing set are shown as solid blue diamonds. Figure 3 shows the out-of-sample validation results for Afghanistan, Belgium, Spain, and Nigeria for the MCAR 40% experiment. Overall, the posterior predictive distributions for imputed values of NER result in plausible time series trends for each country, with the majority of the observations of NER that were simulated as missing captured within the 95% intervals. Results for the same example countries are shown in Figure 4 for the MAR 80% experiment. The predictive intervals are much wider overall for the MAR 80% experiment compared to the MCAR 40% experiment. This is especially apparent for Spain, which has many observations of NER and GER in early years and thus had a high proportion of observed values that were simulated as missing. Despite the very large amount of missing data in the experiments using the 80% rate of simulated missingness, we find the MINTS still imputes plausible country-specific trends that reflect our knowledge of how school enrollment rates change over time.
@@ 5.4 Full Enrollment Data Results
We used MINTS to create 40 multiple imputations for the missing country-years in the original school enrollment data set without simulating any additional country-years as missing. The original data set has a large amount of missing data for NER, with about 73.0% of the 10,302 country-years missing. However, we have partial information about the level of school enrollment arising from GER and the hierarchical time series structure of the data for many of these missing country-years. About 48.8% of the 7,524 country-years where NER is missing have observed values of GER. In these country-years, the coarse measure of school enrollment (GER) provides us with partial information about the refined measure of school enrollment (NER). Partial information about each country’s likely level of enrollment and likely trend over time in enrollment can also be obtained from the hierarchical time series structure of the data. For these reasons, we believe multiple imputation of NER is appropriate despite the large amount of missing data.
The prior distributions for MINTS were specified using the data-based algorithm to determine the value of the control parameters. For the global parameters, the data-based algorithm results in the following prior distributions:
$\sigma_{X}^{2}\sim$ $InvGamma(2,\delta_{X}=11.5),$
$\mu_{drift}\sim$ $N(\nu_{drift}=0.994,\zeta_{drift}^{2}=0.00204),$
$\sigma_{drift^{2}}\sim$ $InvGamma(2,\delta_{drift}=1.87),$
$\sigma_{Y}^{2}\sim$ $InvGamma(2,\delta_{Y}=5.06),$
$\mu_{0}\sim$ $N(0,\zeta_{0}^{2}=0.00233),$
$\sigma_{0}^{2}\sim$ $InvGamma(2,\delta_{0}=2.65).$
The joint prior distribution of $(Y_{c,0},X_{c,0})$ was estimated using an early subset of data that covered years 1970 to 1980 to balance between only including observations close to the first time period and having a sufficient number of observations in the chosen subset. The truncation for the joint prior distribution is specified based on substantive knowledge that $\mathbf{X}$ and $\mathbf{Y}$ are generally increasing over time. The joint prior distribution of $(Y_{c,0},X_{c,0})$ is
[…]
where the truncation is such that $X_{c,0}$ is bounded from below by 0 and from above by the maximum observed value of GER for country $c$. $Y_{c,0}$ is bounded from below by 0 and from above by the minimum of $X_{c,0}$ and the maximum observed value of NER for country $c$.
To check that these prior distributions are appropriate for the school enrollment application, we conducted a prior predictive check and visually compared the prior and posterior distributions for the parameters specified using the data-based algorithm. We found that the chosen prior distributions are sufficiently diffuse and do not overwhelm the posterior. Details of these assessments can be found in Section 9 of the Supplementary Material (Liu and Raftery (2025)).
Figure 5 illustrates the multiply imputed values for NER from MINTS for Afghanistan, Belgium, Spain, and Nigeria. The multiple imputations for NER are shown as translucent grey circles, while the observed values of NER and GER from the original data set are shown as open and solid black circles, respectively. Despite the large amount of missing data for NER in the full enrollment data set, we find MINTS produces imputations that result in plausible imputed trends over time for all countries.
@@
Figure 5: Results of $M=40$ imputations for NER for selected countries. Open and solid black circles indicate observed values of NER and GER, respectively, from the enrollment data set. Translucent grey circles indicate imputed values of NER, where a total of 40 imputations were created for each missing value.
A CSV file containing the 40 completed data sets for both NER and GER and a CSV file containing the median imputed values for both NER and GER are provided in the Supplementary Material (Liu and Raftery (2025)). We note that the intended use of MINTS is to create multiple imputations for the variable of interest NER, and the auxiliary variable GER is only imputed out of necessity. However, we make the imputed values of GER available to provide additional context for how the MINTS method works for the school enrollment application.
We also created multiply imputed estimates for the linear regression of TFR on NER, where $M=40$ multiple imputations were created using MINTS and the existing multiple imputation methods considered in the validation exercises. We estimated the same linear regression of TFR on NER that was used in the analysis model validation in Section 5.2.1, where the quantity of interest is the regression coefficient $\beta_{1}$. Table 4 summarizes the multiply imputed estimates of $\beta_{1}$ from MINTS and the six existing multiple imputation methods using the full enrollment data along with estimates from the data augmentation approach.
In the analysis model validation exercise using the school enrollment data, we found that MINTS resulted in the smallest MAE and closest to nominal coverage, and that pan Fixed Effects was the most competitive alternative out of the existing multiple imputation methods. When applied to the full school enrollment data, MINTS and pan Fixed Effects result in similar multiply imputed estimates of $\beta_{1}$. Both methods estimate that an increase in NER from 0% to 100% is associated with a reduction in TFR of over 4.2 children per woman, where MINTS estimates a smaller effect size than pan Fixed Effects. MINTS and pan Fixed Effects also result in confidence intervals of approximately the same width, but pan Fixed Effects has a smaller estimated FMI.
Table 4: Comparison of multiply imputed estimates of $Q=\beta_{1}$, the regression coefficient on NER in the linear regression of TFR on NER, created using $M=40$ multiple imputations from MINTS and six existing multiple imputation methods. $\bar{Q}_{40}$ denotes the pooled point estimates, 95% CI denotes the 95% confidence intervals for $\bar{Q}_{40}$, and FMI denotes the fraction of missing information as a percentage. $\bar{Q}_{40}$ and the confidence interval bounds are multiplied by 100 before reporting.
Method $\bar{Q}_{40}$ 95% CI FMI MICE PMM -2.24 (-2.40, -2.07) 41.5 pan Fixed -4.36 (-4.52, -4.21) 39.7 pan Random -3.50 (-3.86, -3.14) 90.1 Amelia TS -2.85 (-3.02, -2.68) 34.0 Amelia CS -3.48 (-3.67, -3.29) 58.3 Amelia TSCS -3.55 (-3.79, -3.31) 74.0 MINTS -4.26 (-4.42, -4.11) 53.2 MINTS DA -5.40 (-5.73, -4.84)
The other five models based on existing methods (MICE PMM, pan Random Effects, Amelia TS, Amelia CS, and Amelia TSCS) had substantial bias and undercoverage in the analysis model validation exercise for estimation of $\beta_{1}$. When applied to the full school enrollment data, these five models all lead to multiply imputed point estimates for $\beta_{1}$ that suggest an increase in NER from 0% to 100% is associated with a reduction in TFR of less than 3.6 children per woman. Based on the findings of the analysis model validation exercise, these estimates are likely to be biased compared to the estimates from MINTS and pan Fixed effects and may underestimate the relationship between TFR and NER.
The data augmentation approach based on MINTS results in the largest estimated effect size, with a point estimate for $\beta_{1}$ that suggests an increase in NER from 0% to 100% is associated with a reduction in TFR of about 5.4 children per woman. In the analysis model validation exercise for estimation of $\beta_{1}$, we found that the data augmentation approach had larger MAE than MINTS when the MAR assumption was satisfied. However, the data augmentation approach had the smallest MAE and closest to nominal coverage in the MNAR experiments. If we have reason to suspect a strong violation of the MAR assumption for the full school enrollment data set, then the estimates from MINTS and pan Fixed Effects may also underestimate the true relationship between TFR and NER.
An important limitation of the multiple imputation results presented in this section is that the complete cases in the school enrollment data set only comprise 26.8% of the possible country-years. The MINTS method therefore relies on a relatively small proportion of the data to estimate the relationship between NER and GER. Our validation exercises using the school enrollment data suggest that MINTS produces unbiased multiply imputed estimates with reasonable coverage even in settings with a very high percentage of missing data. However, the small number of complete cases increases the risk that the assumptions of MINTS are violated. The restriction of MINTS to the bivariate setting increases the risk that the ignorability assumption is violated, as the probability of NER being missing for the country-years where both NER and GER are missing may depend on external factors that are not accounted for in the bivariate imputation model. The country-years with both NER and GER missing could also be systematically different from the country-years where at least one measure of school enrollment is observed, and the relationships between NER and GER estimated using the observed country-years may not be appropriate for country-years where both measures of enrollment are missing. If the true missing data mechanism is strongly MNAR, the results from our simulation studies and validation exercises suggest that multiply imputed point estimates from MINTS are likely to be biased and interval estimates are likely to suffer from undercoverage. Incorporation of additional variables that are predictive of the probability of observing NER and GER into the imputation model could improve the plausibility of the MAR assumption, and generalization of the MINTS method to the multivariate setting to accommodate such variables is of interest for future work.